# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v1

@@ 5 Algorithms and Main Results
In this section, we begin by proving a lower regret bound under uniform rewards (see Subsection 5.1), explicitly showing the dependence on the utility for the outside option $v_{0}$. We then introduce OFU-MNL+, an algorithm that achieves minimax optimality, up to logarithmic factors, in scenarios with uniform rewards (Subsection 5.2). Notably, OFU-MNL+ is designed for efficiency, requiring only an $\mathcal{O}(Kd^{3})$ computation cost per iteration and an $\mathcal{O}(d^{2})$ storage cost. Finally, we establish both lower and upper regret bounds under non-uniform rewards (Subsection 5.3).
@@ 5.1 Regret Lower Bound under Uniform Rewards
Discussion of Theorem 1. If $v_{0}=\Theta(1)$, Theorem 1 demonstrates $\Omega(d\sqrt{T/K})$ regret lower bound, which is improved by a factor of $\sqrt{K}$ compared to the one proposed by  Chen et al. [9]  [^2] Note that Chen et al. [9] also consider the uniform reward setting in their proof of the lower bound.. This improvement is mainly attributed to the establishment of a tighter upper bound for the KL divergence (Lemma C.2). Therefore, this clearly indicates that, under uniform rewards, an increase in the assortment size $K$ can lead to an improvement in regret. On the other hand, Chen and Wang [8], Agrawal et al. [5] established regret lower bounds of $\Omega(\sqrt{NT})$ and $\Omega(\sqrt{NT/K})$, respectively, in non-contextual MNL bandits with uniform rewards. They set $v_{0}=\Theta(K)$ to achieve such regrets. Theorem 1 shows that if $v_{0}=\Theta(K)$, we can obtain a regret lower bound of $\Omega(d\sqrt{T})$, which is consistent with the $K$-independent regret in Chen and Wang [8]. To the best of our knowledge, this result is the first to explicitly show the dependency of regret on the utility for the outside option $v_{0}$.
@@ 5.2 Minimax Optimal Regret Upper Bound under Uniform Rewards
Online parameter estimation. Instead of performing MLE as in previous works works [9, 22, 15], inspired by Zhang and Sugiyama [29], we use the mirror descent algorithm to estimate parameter. We first define the multinomial logistic loss function at round $t$ as:
@@
Then, we estimate parameter $\mathbf{w}^{\star}$ as follows:
$\mathbf{w}_{t+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\langle\nabla\ell_{t}(\mathbf{w}_{t}),\mathbf{w}\rangle+\frac{1}{2\eta}\|\mathbf{w}-\mathbf{w}_{t}\|_{\tilde{H}_{t}}^{2}\,,\quad\forall t\geq 1$ (3)
@@
and $H_{t}:=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Note that $\mathcal{G}_{t}(\mathbf{w})=\nabla^{2}\ell_{t}(\mathbf{w})$. This online estimator is efficient in terms of both computation and storage. By a standard online mirror descent formulation [23], (3) can be solved using a single projected gradient step through the following equivalent formula:
@@
Armed with the online estimator, we then proceed to construct the computationally efficient optimistic revenue.
@@
At each time $t$, given the confidence set in Lemma 1, we first calculate the optimistic utility $\alpha_{ti}$ for each item as follows:
$\alpha_{ti}:=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}.$ (5)
@@
$\tilde{R}_{t}(S):=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{ti})},$ (6)
@@
Discussion of Theorem 2. If $T\geq\mathcal{O}(d^{2}(v_{0}+K)^{2}/(\kappa^{2}v_{0}K))$, Theorem 2 shows that our algorithm OFU-MNL+ achieves minimax optimal regret (up to logarithmic factor) in terms of $d$, $T$, $K$, and even $v_{0}$. To the best of our knowledge, ignoring logarithmic factors, our proposed OFU-MNL+ is the first computationally efficient, minimax optimal algorithm in (adversarial) contextual MNL bandits. When $v_{0}=\Theta(1)$, which is the convention in existing MNL bandit literature [21, 22, 15, 3, 29]OFU-MNL+ obtains $\tilde{\mathcal{O}}(d\sqrt{T/K})$ regret. To the best of our knowledge, this is the first algorithm in MNL bandits that shows that as $K$ increases, the regret decreases. And it improves upon the previous regret upper bound of Goyal and Perivier [15] 333 Goyal and Perivier [15] also consider the uniform rewards with $v_{0}=1$., which is $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$, by a factor of $K$. To the best of our knowledge, this is the first algorithm in MNL bandits that shows that as $K$ increases, the regret decreases. And it improves upon the previous regret upper bound of Goyal and Perivier [15], which is $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$, by a factor of $K$. This improvement can be attributed to two key factors: an improved, constant, self-concordant-like property of the loss function (Proposition B.1) and a $K$-free elliptical potential (Lemma D.2). Moreover, by employing an improved bound for the second derivative of the revenue (Lemma D.3), we achieve an enhancement in the regret for the second term, $d^{2}/\kappa$, by a factor of $K^{4}$, in comparison to Goyal and Perivier [15].
When reduced to the logistic bandit, i.e., $K=1$, $r_{t1}=1$, and $v_{0}=1$, our algorithm can also achieve a regret of $\tilde{\mathcal{O}}(d\sqrt{\kappa T})$ by Corollary 1 in Zhang and Sugiyama [29], which is consistent with the results in Abeille et al. [2], Faury et al. [13].
@@
Since we employ online parameter estimation, akin to Zhang and Sugiyama [29], our algorithm demonstrates computational efficiency in parameter estimation, incurring only incurring $\mathcal{O}(Kd^{3})$ computation cost and $\mathcal{O}(d^{2})$ storage cost, which are completely independent of $t$. Furthermore, a naive approach to selecting the optimistic assortment requires enumerating all possible ($N$ choose $K$) assortments, resulting in exponential computational cost [9]. However, by constructing the optimistic expected revenue according to (6) (inspired by Oh and Iyengar [22]), our algorithm needs only $\mathcal{O}(N)$ computational cost.
@@ 5.3 Regret Upper & Lower Bounds under Non-Uniform Rewards
Discussion of Theorem 3. In contrast to Theorem 1, which considers uniform rewards, the regret lower bound is independent of the assortment size $K$. Note that Theorem 3 does not claim that non-uniform rewards inherently make the problem more difficult. Rather, it implies that there exists an instance with adversarial non-uniform rewards, where regret does not improve even with an increase in $K$. Moreover, the assumption that $v_{0}=\Theta(1)$ is natural and common in the existing literature on contextual MNL bandits [21, 22, 15, 3, 29] (refer Appendix B.1). To the best of our knowledge, this is the first established lower bound for non-uniform rewards in MNL bandits.
We also prove a nearly matching upper bound. The algorithm OFU-MNL+ is also applicable in the case of non-uniform rewards. However, because the optimistic expected revenue $\tilde{R}_{t}(S)$ is no longer an increasing function of $\alpha_{ti}$, optimizing for $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$ no longer equates to simply selecting the top $K$ items with the highest optimistic utility. Instead, we employ assortment optimization methods introduced in Rusmevichientong et al. [25], Davis et al. [12], which are efficient polynomial-time (independent of $t$ [^4] An interior point method would generally solve the problem with a computational complexity of $\mathcal{O}(N^{3.5})$. algorithms available for solving this optimization problem. Therefore, our algorithm OFU-MNL+ is also computationally efficient under non-uniform rewards.
@@
Under the same assumptions and parameter settings as Theorem 2, if the rewards are non-uniform and $v_{0}=\Theta(1)$, then with a probability of at least $1-\delta$, the cumulative regret of OFU-MNL+ is upper-bounded by
@@
Discussion of Theorem 4. If $T\geq\mathcal{O}(d^{2}/\kappa^{2})$, our algorithm achieves a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$ when the reward for each item is non-uniform. This performance is minimax optimal up to a logarithmic factor. The rewards assumption in Zhang and Sugiyama [29], which posits that the norm of the reward vector $\bm{\rho}_{t}=[r_{t1},\dots r_{t|S_{t}|}]^{\top}\in\mathbb{R}^{|S_{t}|}$ is bounded by a constant, is stronger than our assumption. We allow the norm’s bound to scale with $K$. Consequently, we need a novel approach to achieve regret that does not scale with $K$. We centralize the features and propose the novel elliptical potential lemma for them, as detailed in Lemma G.2. Furthermore, our algorithm is capable of achieving $1/\kappa$-free regret (in a leading term) under both uniform and non-uniform rewards. In contrast, the algorithm in Goyal and Perivier [15] is limited to achieving this only in the uniform reward setting.
@@ 6 Numerical Experiments
In this section, we empirically evaluate the performance of our algorithm OFU-MNL+. We measure cumulative regret over $T=3000$ rounds. For each experimental setup, we run the algorithms across $20$ independent instances and report the average performance. In each instance, the underlying parameter $\mathbf{w}^{\star}$ is randomly sampled from a $d$-dimensional uniform distribution, where each element of $\mathbf{w}^{\star}$ lies within the range $[-1/\sqrt{d},1/\sqrt{d}]$ and is not known to the algorithms. Additionally, the context features $x_{ti}$ are drawn from a $d$-dimensional multivariate Gaussian distribution, with each element of $x_{ti}$ clipped to the range $[-1/\sqrt{d},1/\sqrt{d}]$. This setup ensures compliance with Assumption 1. In the uniform reward setting (first row of Figure 1), the combinatorial optimization step to choose the assortment reduces to sorting items by their utility estimate. In the non-uniform reward setting (second row of Figure 1), the rewards are sampled from a uniform distribution in each round, i.e., $r_{ti}\sim\operatorname{Unif}(0,1)$. Refer Appendix H for more details.
We compare the performance of OFU-MNL+ with those of the practical and state-of-the-art algorithms: the Upper Confidence Bound-based algorithm, UCB-MNL [21], and the Thompson Sampling-based algorithm, TS-MNL [21]. Figure 1 demonstrates that our algorithm significantly outperforms other baseline algorithms. In the uniform reward setting, as $K$ increases, the cumulative regrets of all algorithms tend to decrease. In contrast, this trend is not observed in the non-uniform reward setting. Furthermore, the results also show that our algorithm maintains a constant computation cost per round, while the other algorithms exhibit a linear dependence on $t$. In Appendix H, we present the additional running time curves (Figure H.1) as well as the regret curves of the other configuration where $v_{0}=\Theta(K)$ (Figure H.2). All of these empirical results align with our theoretical results.

### v6

@@ 5 Algorithms and Main Results
In this section, we begin by proving the tightest regret lower bound under uniform rewards (Subsection 5.1), explicitly showing the dependence on the attraction parameter for the outside option $v_{0}$. We then introduce OFU-MNL+, an algorithm that achieves minimax optimality, up to logarithmic factors, under uniform rewards (Subsection 5.2). Notably, OFU-MNL+ is designed for efficiency, requiring only constant computation and storage costs. Finally, we establish the tightest regret lower bound and a matching minimax optimal regret upper bound under non-uniform rewards (Subsection 5.3).
@@ 5.1 Regret Lower Bound under Uniform Rewards
[…]
@@ 5.2 Minimax Optimal Regret Upper Bound under Uniform Rewards
Online parameter estimation. Instead of performing MLE as in previous works [14, 40, 43], inspired by Zhang and Sugiyama [52], we use the online mirror descent algorithm to estimate the parameter. We first define the multinomial logistic loss function at round $t$ as:
@@
In Proposition C.1, we will show that the loss function has the constant parameter self-concordant-like property. We estimate the true parameter $\mathbf{w}^{\star}$ as follows:
$\mathbf{w}_{t+1}=\mathop{\mathrm{argmin}}_{\mathbf{w}\in\mathcal{W}}\langle\nabla\ell_{t}(\mathbf{w}_{t}),\mathbf{w}\rangle+\frac{1}{2\eta}\|\mathbf{w}-\mathbf{w}_{t}\|_{\tilde{H}_{t}}^{2}\,,\quad\forall t\geq 1,$ (3)
@@
and $H_{t}:=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Note that $\mathcal{G}_{t}(\mathbf{w})=\nabla^{2}\ell_{t}(\mathbf{w})$. This online estimator is efficient in terms of both computation and storage. By a standard online mirror descent formulation [41], the optimization problem in (3) can be solved using a single projected gradient step through the following equivalent formula:
@@
Armed with the online estimator, we construct the computationally efficient optimistic revenue.
@@
At each time $t$, given the confidence set in Lemma 1, we first calculate the optimistic utility $\alpha_{ti}$ as:
$\alpha_{ti}:=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}},\quad\,\forall i\in[N].$ (5)
@@
$\tilde{R}_{t}(S):=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})},$ (6)
@@
Discussion of Theorem 2. If $T\geq\mathcal{O}(d^{2}(v_{0}+K)^{2}/(\kappa^{2}v_{0}K))$, Theorem 2 shows that our algorithm OFU-MNL+ achieves minimax optimal regret (up to logarithmic factor) in terms of all $d$, $T$, $K$, and even $v_{0}$. To the best of our knowledge, ignoring logarithmic factors, our proposed algorithm is the first computationally efficient, minimax optimal algorithm in (adversarial) contextual MNL bandits. When $v_{0}=\Theta(1)$, which is the convention in existing MNL bandit literature [39, 40, 43, 4, 52]OFU-MNL+ obtains $\tilde{\mathcal{O}}(d\sqrt{T/K})$ regret. This represents an improvement over the previous upper bound of Perivier and Goyal [43] 222 Perivier and Goyal [43] also consider the uniform rewards ($r_{ti}=1$) with $v_{0}=1$., which is $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$, by a factor of $K$. This improvement can be attributed to two key factors: an improved, constant, self-concordant-like property of the loss function (Proposition C.1) and a $K$-free elliptical potential lemma (Lemma E.2). Furthermore, by employing an improved bound for the second derivative of the revenue (Lemma E.3), we achieve an enhancement in the regret for the second term, $d^{2}/\kappa$, by a factor of $K^{4}$, in comparison to Perivier and Goyal [43]. Unless $v_{0}=\Theta(K)$, Theorem 2 indicates that the regret decreases as the assortment size $K$ increases. To the best of our knowledge, this is the first algorithm in MNL bandits to show that increasing $K$ results in a reduction in regret. Moreover, when reduced to the logistic bandit, i.e., $K=1$, $r_{t1}=1$, and $v_{0}=1$, our algorithm can also achieve a regret of $\tilde{\mathcal{O}}(d\sqrt{\kappa T})$ by Corollary 1 in Zhang and Sugiyama [52], which is consistent with the results in Abeille et al. [3], Faury et al. [21]. The proof is deferred to Appendix E.
@@
The proposed algorithm is computationally efficient in both parameter updates and assortment selections. Since we employ online parameter estimation, akin to Zhang and Sugiyama [52], our algorithm demonstrates computational efficiency in parameter estimation, incurring only incurring $\mathcal{O}(Kd^{3})$ computation cost and $\mathcal{O}(d^{2})$ storage cost, which are completely independent of $t$. Furthermore, a naive approach to selecting the optimistic assortment requires enumerating all possible ($N$ choose $K$) assortments, resulting in exponential computational cost [14]. However, by constructing the optimistic expected revenue according to (6) (inspired by Oh and Iyengar [40]), our algorithm needs only $\mathcal{O}(N)$ computational cost.
@@ 5.3 Regret Upper & Lower Bounds under Non-Uniform Rewards
Discussion of Theorem 3. In contrast to Theorem 1, which considers uniform rewards, the regret lower bound is independent of the assortment size $K$. Note that Theorem 3 does not claim that non-uniform rewards inherently make the problem more difficult. Rather, it implies that there exists an instance with adversarial non-uniform rewards, where regret does not improve even with an increase in $K$. Moreover, the assumption that $v_{0}=\Theta(1)$ is common in the existing literature on contextual MNL bandits [39, 40, 43, 4, 52] (refer Appendix C.1). To the best of our knowledge, this is the first established lower bound for non-uniform rewards in MNL bandits, even in the non-contextual setting. The proof is deferred to Appendix G.
We also prove a matching upper bound up to logarithmic factors. The algorithm OFU-MNL+ is also applicable in the case of non-uniform rewards. However, because the optimistic expected revenue $\tilde{R}_{t}(S)$ is no longer an increasing function of $\alpha_{ti}$, optimizing for $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$ no longer equates to simply selecting the top $K$ items with the highest optimistic utility. Instead, we employ assortment optimization methods introduced in Rusmevichientong et al. [46], Davis et al. [19], which are efficient polynomial-time (independent of $t$ [^3] An interior point method would generally solve the problem with a computational complexity of $\mathcal{O}(N^{3.5})$. algorithms available for solving this optimization problem. Therefore, our algorithm is also computationally efficient under non-uniform rewards.
@@
Under the same assumptions and parameter settings as Theorem 2, if the rewards are non-uniform and $v_{0}=\Theta(1)$, then with a probability at least $1-\delta$, the cumulative regret of OFU-MNL+ is upper-bounded by
@@
Discussion of Theorem 4. If $T\geq\mathcal{O}(d^{2}/\kappa^{2})$, our algorithm achieves a regret of $\tilde{\mathcal{O}}(d\sqrt{T})$ when the reward for each item is non-uniform, demonstrating that OFU-MNL+ is minimax optimal up to a logarithmic factor. Recall that we relax the bounded assumption on the reward compared to Zhang and Sugiyama [52] (refer Assumption 1); thus, we allow the sum of the squared rewards in the assortment to scale with $K$. Consequently, we need a novel approach to achieve the regret that does not scale with $K$. To this end, we centralize the features, i.e., $\tilde{x}_{ti}=x_{ti}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}_{t+1})}[x_{tj}]$, and propose a novel elliptical potential lemma for them, as detailed in Lemma H.3. Note that our algorithm is capable of achieving $1/\kappa$-free regret (in the leading term) under both uniform and non-uniform rewards. In contrast, the algorithm in Perivier and Goyal [43] is limited to achieving this only in the uniform reward setting. Furthermore, compared to the regret bound in Chen et al. [14], which is $\tilde{\mathcal{O}}(d\sqrt{T})$, our regret bounds has the same order of regret with theirs. However, their algorithm is computationally intractable as it requires enumerating all possible assortments, whereas our algorithm incurs only a constant computational cost per round. The proof is deferred to Appendix H.
6 Instance-Dependent Bounds
In this section, we show that instance-dependent upper and lower bounds are also achievable under uniform rewards. We define the degree of non-linearity for the optimal assortment $S^{\star}_{t}$ at round $t$ under the true parameter $\mathbf{w}^{\star}$ as $\kappa^{\star}_{t}:=\sum_{i\in S^{\star}_{t}}p_{t}(i|S^{\star}_{t},\mathbf{w}^{\star})p_{t}(0|S^{\star}_{t},\mathbf{w}^{\star})$. We first establish the instance-dependent lower bound under uniform rewards.
Proposition 1 (Instance-dependent regret lower bound, Uniform rewards).
Under the same conditions as Theorem 1, for any policy $\pi$ and for $T\geq d^{2}/\kappa$, there exists a worst-case problem instance such that the worst-case expected regret of $\pi$ is lower bounded as follows:
$\sup_{\mathbf{w}}\mathbb{E}^{\pi}_{\mathbf{w}}\left[\textup{{Reg}}_{T}(\mathbf{w})\right]=\Omega\left(d\sqrt{\sum_{t=1}^{T}\kappa^{\star}_{t}}\right).$
The proof is deferred to Appendix I. We also provide the matching upper bound.
Proposition 2 (Instance-dependent regret upper bound of OFU-MNL+, Uniform rewards).
Under the same assumptions, parameter settings, and reward structure as Theorem 2, with a probability at least $1-\delta$, the cumulative regret of OFU-MNL+ is upper-bounded by
$\textup{{Reg}}_{T}(\mathbf{w}^{\star})$ $=\tilde{\mathcal{O}}\left(d\sqrt{\sum_{t=1}^{T}\kappa^{\star}_{t}}+\frac{1}{\kappa}d^{2}\right).$
The proof is deferred to Appendix J. For sufficiently large $T$, the regret upper bound (Proposition 2) matches the regret lower bound (Proposition 1), up to logarithmic factor. To the best of our knowledge, these are the first minimax instance-dependent regret bounds under uniform rewards. Note that, in the worst case, $\kappa^{\star}_{t}=\tilde{\mathcal{O}}(\sqrt{v_{0}K}/(v_{0}+K))$, which indicates that these results provide a strict improvement over the worst-case bounds given in Theorems 1 and 2.
Some readers may expect instance-dependent regret bounds for non-uniform rewards as well. Unfortunately, we were unable to establish these. Recall that $\kappa^{\star}_{t}$ represents the degree of non-linearity for the optimal assortment $S^{\star}_{t}$. However, in the proofs for regret bounds, we encounter terms associated with the chosen assortment $S_{t}$, such as $\sum_{i\in S^{\star}_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})$. To address this, we use the mean value theorem-based analysis (Lemma I.5) to replace this quantity with $\kappa^{\star}_{t}$. Under non-uniform rewards, however, the mean value theorem does not apply because the sizes and rewards of $S^{\star}_{t}$ and $S_{t}$ may differ. Addressing this problem would be an interesting direction for future research.
@@ 7 Numerical Experiments
In this section, we empirically evaluate the performance of our algorithm OFU-MNL+. We measure cumulative regret over $T=3000$ rounds. For each experimental setup, we run the algorithms across $20$ independent instances and report the average performance. In each instance, the underlying parameter $\mathbf{w}^{\star}$ is randomly sampled from a $d$-dimensional uniform distribution, where each element of $\mathbf{w}^{\star}$ lies within the range $[-1/\sqrt{d},1/\sqrt{d}]$ and is not known to the algorithms. Additionally, the context features $x_{ti}$ are drawn from a $d$-dimensional multivariate Gaussian distribution, with each element of $x_{ti}$ clipped to the range $[-1/\sqrt{d},1/\sqrt{d}]$. This setup ensures compliance with Assumption 1. In the uniform reward setting (first row of Figure 1), the combinatorial optimization step to choose the assortment reduces to sorting items by their utility estimate. In the non-uniform reward setting (second row of Figure 1), the rewards are sampled from a uniform distribution in each round, i.e., $r_{ti}\sim\operatorname{Unif}(0,1)$. Refer Appendix K for more details.
We compare the performance of OFU-MNL+ with those of the practical and state-of-the-art algorithms: the Upper Confidence Bound-based algorithm, UCB-MNL [39], and the Thompson Sampling-based algorithm, TS-MNL [39]. Figure 1 demonstrates that our algorithm significantly outperforms other baseline algorithms. In the uniform reward setting, as $K$ increases, the cumulative regrets of all algorithms tend to decrease. In contrast, this trend is not observed in the non-uniform reward setting. Furthermore, the results also show that our algorithm maintains a constant computation cost per round, while the other algorithms exhibit a linear dependence on $t$. In Appendix K, we present the additional runtime curves (Figure K.1) as well as the regret curves of the other configuration where $v_{0}=\Theta(K)$ (Figure K.2). All of these empirical results align with our theoretical results.